# Self-Organization and Spectral Mechanism of Attractor Landscapes in High-Capacity Kernel Hopfield Networks

### v3

@@ 2.3 Geometry and Spectrum of Neural Dynamics
Understanding the geometry of neural dynamics has been a longstanding goal, often approached through statistical mechanics [14] or information geometry [15]. Amari and Wu analyzed the spatial distribution of critical points in neural fields, providing insights into pattern stability. Our work complements these theoretical foundations by introducing a data-driven spectral perspective. While statistical mechanics often assumes the thermodynamic limit ($N\to\infty$) with random patterns, our spectral analysis of the “Ridge” and “Spectral Concentration” provides a concrete mechanism for finite-size networks. It explains how the system self-organizes its weight spectrum to balance stability (via the leading eigenvalue) and diversity (via the trailing eigenvalues), offering a bridge between abstract capacity bounds and the actual geometric properties of learned attractors.
@@ 4.2 Decomposing the Forces: The Mechanism of Self-Organization
This strong anti-correlation provides the key to understanding the ridge. The total sharpness, $\|\nabla V\|^{2}$, is given by $\|\bm{F}_{d}\|^{2}+\|\bm{F}_{i}\|^{2}+2\left<\bm{F}_{d},\bm{F}_{i}\right$>. When $\rho\approx-1$, this simplifies to $\|\nabla V\|^{2}\approx(\|\bm{F}_{d}\|-\|\bm{F}_{i}\|)^{2}$. This means that the total sharpness is determined by the difference in the magnitudes of the two opposing forces.
@@ 4.3 The Nature of the Optimization Ridge
Figure 3: rowth of force component magnitudes as a function of storage load $P/N$ for a fixed small $\gamma$. The Direct Force (blue, $\|\bm{F}_{d}\|^{2}$) grows significantly faster than the Indirect Force (green, $\|\bm{F}_{i}\|^{2}$). Theoretically, this reflects the amplification of the leading spectral mode ($\lambda_{1}$) due to Spectral Concentration, which allows the driving force to overwhelm the interference and create a deep attractor basin.
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
On the Ridge of Optimization, the matrix $\bm{\alpha}$ exhibits strong spectral concentration. The leading eigenvalue becomes dominant ($\lambda_{1}\gg\lambda_{k}$ for $k>1$), thereby maximizing the Pinnacle Sharpness via Proposition 1. Crucially, the trailing eigenvalues remain non-zero ($\lambda_{k}>0$), maintaining the full-rank structure necessary to store multiple patterns.
@@ 6.3 Practical Implications and Future Directions
The clean, spurious-free landscape achieved by Spectral Concentration makes KLR Hopfield networks an attractive candidate for reliable content-addressable memory systems. However, the computational cost of kernel evaluations scales with the number of stored patterns. Future work should explore sparse approximations (e.g., Nyström method or Random Fourier Features) to reduce the effective rank explicitly, potentially enforcing the “Spectral Concentration” structure by design to achieve linear scaling. Furthermore, extending this spectral theory to analyze the dynamics of Transformer attention layers, which share mathematical similarities with kernel memories, remains an exciting avenue for research.
Furthermore, our discovery of Spectral Concentration provides a theoretical justification for using low-rank approximations in kernel machines. The fact that the effective weight matrix naturally exhibits a low stable rank on the Ridge implies that sparse approximation methods, such as the Nyström method [16], can capture the essential dynamics without significant loss of performance, paving the way for scalable implementations.
@@ 7 Conclusion
Our phenomenological analysis identified a “Ridge of Optimization” in the phase diagram, where the network achieves maximal stability characterized by a unique “Force Antagonism.” Crucially, our theoretical analysis revealed the spectral origin of this phenomenon: Spectral Concentration. We demonstrated that on the Ridge, the learning algorithm drives the weight spectrum into a critical state—dominated by a massive leading eigenvalue for global stability, yet retaining a non-zero heavy tail for high-capacity storage. This finding dispels the notion of a simple rank-1 collapse and instead portrays the optimal memory landscape as a finely tuned balance between attractive force and pattern discriminability.
@@ A.2 Proof of Proposition 1 (Stability-Spectrum Link)
We aim to show that the Hessian of the energy function, which governs stability, is spectrally related to the Gram matrix.
The energy function is $E(\bm{s})=-\frac{1}{2}\langle\phi(\bm{s}),\bm{W}\phi(\bm{s})\rangle$. Let $\bm{s}^{*}$ be a fixed point. The Hessian $H=\nabla^{2}E(\bm{s}^{*})$ is given by $H=J^{\top}\bm{W}J$, where $J$ is the Jacobian of the feature map $\phi(\bm{s})$ at $\bm{s}^{*}$. Substituting $\bm{W}=\bm{\Phi}\bm{\alpha}\bm{\Phi}^{\top}$, we have:
$H=J^{\top}\bm{\Phi}\bm{\alpha}\bm{\Phi}^{\top}J=(\bm{\Phi}^{\top}J)^{\top}\bm{\alpha}(\bm{\Phi}^{\top}J).$ (11)
Justification of the Isometry Approximation: To relate the spectrum of $H$ to the kernel matrix $\bm{K}=\bm{\Phi}^{\top}\bm{\Phi}$, we analyze the term $\bm{A}=\bm{\Phi}^{\top}J$. The matrix $\bm{A}\bm{A}^{\top}=\bm{\Phi}^{\top}(JJ^{\top})\bm{\Phi}$ represents the projection of the feature-space Jacobian onto the subspace spanned by the patterns. For high-dimensional random patterns and radial basis function (RBF) kernels, the feature mapping $\phi(\cdot)$ behaves locally as an approximate isometry on the tangent space of the data manifold. Under this condition, the operator $JJ^{\top}$ acts approximately as a scalar scaling factor on the relevant subspace:
$JJ^{\top}\approx c\cdot\bm{I}_{\mathcal{H}}\quad\text{(on the span of patterns)},$ (12)
where $c>0$ is a constant related to the kernel width and input dimension. Using this approximation, we obtain:
$\bm{A}\bm{A}^{\top}=\bm{\Phi}^{\top}(c\bm{I})\bm{\Phi}=c\bm{K}.$ (13)
Consequently, the non-zero eigenvalues of $H$ are proportional to the non-zero eigenvalues of the matrix:
$\bm{M}=\bm{\alpha}^{1/2}\bm{K}\bm{\alpha}^{1/2}.$ (14)
Therefore, the Pinnacle Sharpness $M(\bm{\xi}^{\mu})$, which corresponds to the squared gradient norm (and relates to the dominant curvature), scales with the largest eigenvalue:
$M\propto\lambda_{\max}(\bm{\alpha}\bm{K}).$ (15)

### v4

@@ 2.3 Geometry and Spectrum of Neural Dynamics
Understanding the geometry of neural dynamics has been a longstanding goal, often approached through statistical mechanics [14] or information geometry [15]. Amari and Wu pioneered the geometric analysis of kernel machines, demonstrating how a kernel function induces a Riemannian metric in the input space that governs the spatial resolution of classification [15]. Our work complements these theoretical foundations by introducing a data-driven spectral perspective. While statistical mechanics often assumes the thermodynamic limit ($N\to\infty$) with random patterns, and classical information geometry focuses on static classifiers, our spectral analysis of the “Ridge” and “Spectral Concentration” provides a concrete mechanism for the dynamics of finite-size networks. It explains how the system self-organizes its weight spectrum to balance stability (via the leading eigenvalue) and diversity (via the trailing eigenvalues), offering a bridge between abstract geometric theories and the actual stability of learned attractors.
@@ 4.2 Decomposing the Forces: The Mechanism of Self-Organization
This strong anti-correlation provides the key to understanding the ridge. The Pinnacle Sharpness, $\|\nabla V\|^{2}$, is given by $\|\bm{F}_{d}\|^{2}+\|\bm{F}_{i}\|^{2}+2\left<\bm{F}_{d},\bm{F}_{i}\right$>. When $\rho\approx-1$, this simplifies to $\|\nabla V\|^{2}\approx(\|\bm{F}_{d}\|-\|\bm{F}_{i}\|)^{2}$. This means that the Pinnacle Sharpness is determined by the difference in the magnitudes of the two opposing forces.
@@ 4.3 The Nature of the Optimization Ridge
Figure 3: Growth of force component magnitudes as a function of storage load $P/N$ for a fixed small $\gamma$. The Direct Force (blue, $\|\bm{F}_{d}\|^{2}$) grows significantly faster than the Indirect Force (green, $\|\bm{F}_{i}\|^{2}$). Theoretically, this reflects the amplification of the leading spectral mode ($\lambda_{1}$) due to Spectral Concentration, which allows the driving force to overwhelm the interference and create a deep attractor basin.
@@ 5.1.2 Step 2: Spectral Concentration on the Ridge
On the Ridge of Optimization, the matrix $\bm{\alpha}$ exhibits strong spectral concentration. The leading eigenvalue becomes dominant ($\lambda_{1}\gg\lambda_{k}$ for $k>1$), thereby maximizing the Pinnacle Sharpness via Proposition 1. Crucially, unlike a rank-1 collapse, the trailing eigenvalues remain non-zero ($\lambda_{k}>0$), maintaining the full-rank structure necessary to store multiple patterns.
Appendix A.3 provides a theoretical justification for this phenomenon. It explains how the maximum-margin objective naturally leads to a spectral structure that amplifies the principal mode for stability while preserving separability for capacity.
@@ 6.3 Practical Implications and Future Directions
The clean, spurious-free landscape achieved by Spectral Concentration makes KLR Hopfield networks an attractive candidate for reliable content-addressable memory systems. However, the computational cost of kernel evaluations scales with the number of stored patterns. Future work should explore sparse approximations (e.g., Nyström method [16] or Random Fourier Features [17]) to reduce the effective rank explicitly, potentially enforcing the “Spectral Concentration” structure by design to achieve linear scaling. Furthermore, extending this spectral theory to analyze the dynamics of Transformer attention layers, which share mathematical similarities with kernel memories, remains an exciting avenue for research.
Furthermore, our discovery of Spectral Concentration provides a theoretical justification for using low-rank approximations in kernel machines. The fact that the effective weight matrix naturally exhibits a low stable rank on the Ridge implies that sparse approximation methods, such as the Nyström method, can capture the essential dynamics without significant loss of performance, paving the way for scalable implementations.
@@ 7 Conclusion
Our phenomenological analysis identified a “Ridge of Optimization” in the phase diagram, where the network achieves maximal stability characterized by a unique “Force Antagonism.” Crucially, our theoretical analysis revealed the spectral origin of this phenomenon: Spectral Concentration. We demonstrated that on the Ridge, the learning algorithm drives the weight spectrum into a critical state characterized by a massive leading eigenvalue for global stability, while retaining a non-zero heavy tail for high-capacity storage. This finding dispels the notion of a simple rank-1 collapse and instead portrays the optimal memory landscape as a finely tuned balance between attractive force and pattern discriminability.
@@ A.2 Proof of Proposition 1 (Stability-Spectrum Link)
We aim to show that the Hessian of the energy function, which governs the local stability in the input space, is spectrally related to the kernel Gram matrix in the feature space.
The energy function is defined as $E(\bm{s})=-\frac{1}{2}\langle\phi(\bm{s}),\bm{W}\phi(\bm{s})\rangle$. Let $\bm{s}^{*}$ be a fixed point. The Hessian matrix $H=\nabla^{2}E(\bm{s}^{*})$ is given by:
$H=J^{\top}\bm{W}J,$ (11)
where $J$ is the Jacobian matrix of the feature map $\phi(\bm{s})$ at $\bm{s}^{*}$.
To analyze the spectrum of $H$, consider the quadratic form with respect to an arbitrary unit vector $\bm{u}$ in the input space ($\|\bm{u}\|^{2}=1$):
$\bm{u}^{\top}H\bm{u}=\bm{u}^{\top}(J^{\top}\bm{W}J)\bm{u}=(J\bm{u})^{\top}\bm{W}(J\bm{u}).$ (12)
Let $\bm{v}=J\bm{u}$ be the mapped vector in the feature space. A key property of the RBF kernel is its local isometry; the mapping scales local distances uniformly. Specifically, using the Taylor expansion of the kernel, one can show that the squared norm scales as $\|\bm{v}\|^{2}=\bm{u}^{\top}(J^{\top}J)\bm{u}\approx 2\gamma\|\bm{u}\|^{2}=2\gamma$.
Thus, maximizing the curvature $\bm{u}^{\top}H\bm{u}$ in the input space is equivalent to maximizing the quadratic form $\bm{v}^{\top}\bm{W}\bm{v}$ in the feature space, subject to the norm constraint on $\bm{v}$. Substituting the weight operator $\bm{W}=\bm{\Phi}\bm{\alpha}\bm{\Phi}^{\top}$, the relevant spectral properties are determined by the non-zero eigenvalues of the matrix product $\bm{\Phi}^{\top}\bm{W}\bm{\Phi}$ (restricted to the span of patterns), which reduces to analyzing the effective Gram matrix:
$\bm{K}_{\alpha}=\bm{\alpha}^{1/2}\bm{K}\bm{\alpha}^{1/2}.$ (13)
Note that the non-zero eigenvalues of $\bm{K}_{\alpha}$ are identical to those of $\bm{\alpha}\bm{K}$. Consequently, the Pinnacle Sharpness $M(\bm{\xi}^{\mu})$, which corresponds to the squared gradient norm (and scales with the dominant curvature), is proportional to the leading eigenvalue of this matrix:
$M\propto\lambda_{\max}(\bm{K}_{\alpha}).$ (14)